OpenStudy (anonymous):

Solve for x to the nearest tenth. Sin 34= 5/x

OpenStudy (campbell_st):

so making x the subject. \[x = \frac{5}{\sin(34)}\] now you can enter it into a calculator/

OpenStudy (anonymous):

I got a decimal answer @campbell_st

OpenStudy (campbell_st):

yes that would be correct so if you round your decimal to the 1st decimal place...you'll get your answer. It should be larger than 5...

OpenStudy (anonymous):

i got 14 @campbell_st

OpenStudy (campbell_st):

I got 8.9

OpenStudy (anonymous):

cos 49 = x/9 how would i do that? @campbell_st

OpenStudy (campbell_st):

well make x the subject... so multiply both sides of the equation by 9 then 9*cos(49) = x hope this helps...